OpenStudy (anonymous):

The perimeter of a triangle is 76cm. Side a of the triangle is twice as long as side b. Side c is 1cm longer than side a. Find the length of each side.

OpenStudy (anonymous):

2b + b + 2b +1 = 76 cm 5b = 75 cm b = 15 cm a = 30 cm b = 15 cm c = 31 cm

OpenStudy (anonymous):

How did you find the individual length of each side

OpenStudy (anonymous):

By being clever ;)

OpenStudy (anonymous):

a = 2b = 2 * 15 = 30 b = 15 c = a +1 = 2b+1 = 30 +1 = 31

OpenStudy (anonymous):

Ok i get it Thanks!